OpenStudy (anonymous):

one root of x^ - px + q=0 exceeds the other by 3 then 4q+ 9 is?

OpenStudy (anonymous):

i took the roost as y and y+3

OpenStudy (experimentx):

see what you can do for it

OpenStudy (anonymous):

4q+ 9 = 4y^2 + 12y +9

OpenStudy (anonymous):

b^2 - 4ac =0

OpenStudy (anonymous):

the option are (a) a perfect square (b) a perfect cube

OpenStudy (experimentx):

(2y+3)^2 = ?

OpenStudy (anonymous):

yup... i took the discriminant to see the .. thxxx

OpenStudy (anonymous):

can we take discriminant to see?? @experimentX

OpenStudy (experimentx):

lol ... no need. just see ... what you have in your expression and what are options on your answer.

OpenStudy (anonymous):

ok!...